OpenStudy (anonymous):

help please. classify as arithmetic, geometric, or neither. Finally find an explicit formula (rule for the nth term) for each. Designate the first term with n = 1. ii)2b,4b,6b,8b,10b,12b,14b

OpenStudy (anonymous):

if a2-a1=a3-a2=a4-a3=...=an-an-1 ,then arithmetic. if a2/a1=a3/a2=a4/a3=...=an/an-1,then geometric.

Directrix (directrix):

The common difference is 2b. That is, each term is 2b more than the previous term. So, is the sequence arithmetic or geometric?